OpenStudy (anonymous):

WILL GIVE MEDAL The container that holds the water for the football team is 1/6 full. After pouring in 7 gallons of water, it is 2/3 full. How many gallons can the container hold?

OpenStudy (taylor0402):

18 gallons of water

OpenStudy (anonymous):

18 isn't right :(

jimthompson5910 (jim_thompson5910):

x = capacity of container (in gallons) Starting amount: x/6 amount we pour out: 7 amount we have left: 2x/3 so the equation is (x/6) - 7 = 2x/3 hopefully you can see how I am getting this

OpenStudy (taylor0402):

oops sorry wrong question my bad

jimthompson5910 (jim_thompson5910):

oh I made a bad typo, they are pouring water IN, not out let me fix

jimthompson5910 (jim_thompson5910):

x = capacity of container (in gallons) Starting amount: x/6 amount we pour in: 7 amount we have now: 2x/3 so the equation is (x/6) + 7 = 2x/3

OpenStudy (anonymous):

THANK YOU SO MUCH here's the next question if you guys are down to help I'd be so happy Linda's penny bank is 1/2 full. After she removes 200 pennies, it is 3/10 full. How many pennies can Linda's bank hold?

jimthompson5910 (jim_thompson5910):

`Linda's penny bank is 1/2 full. After she removes 200 pennies, it is 3/10 full. How many pennies can Linda's bank hold? ` x = total number of pennies the container can hold starts off at x/2 pennies remove 200 to be left with (x/2) - 200 we're left with 3x/10 pennies since "it is 3/10 full" after we pull those 200 pennies out so... (x/2) - 200 = 3x/10

OpenStudy (anonymous):

PERFECT I GOT IT!!! I have ONE more question and then I'm done with the lesson, it's this Lashonda's penny bank is 3/8 full. After she adds 240 pennies, it is 3/4 full. How many pennies can Lashonda's bank hold?

jimthompson5910 (jim_thompson5910):

what do you think the equation will be in this case?

OpenStudy (anonymous):

Hm, 3/8 + 240 = 3/4? I know that's not right damn it

jimthompson5910 (jim_thompson5910):

x is involved somehow x = total number of pennies the container can hold

OpenStudy (anonymous):

I'm so bad at word problems unless there's an equation in front of me

jimthompson5910 (jim_thompson5910):

x = total number of pennies the container can hold `Lashonda's penny bank is 3/8 full` how many pennies are currently in the penny bank?

OpenStudy (anonymous):

I'm not sure honestly. I really can't tell you

OpenStudy (anonymous):

I sat there and tried but its not happening

jimthompson5910 (jim_thompson5910):

wouldn't you agree it's simply 3/8 times x?

OpenStudy (anonymous):

Oh yeah, that I could've told you I just thought you meant like whats the actual number haha

jimthompson5910 (jim_thompson5910):

no I want something in terms of x

jimthompson5910 (jim_thompson5910):

so we have 3/8 times x which is the same as 3x/8

jimthompson5910 (jim_thompson5910):

`After she adds 240 pennies, it is 3/4 full` do you know what equation forms?

OpenStudy (anonymous):

3x/8 + 240 = 3/4?

jimthompson5910 (jim_thompson5910):

very close

jimthompson5910 (jim_thompson5910):

it's not going to simply be 3/4 though

OpenStudy (anonymous):

So, 3x/8 + 240= 3/4x?

jimthompson5910 (jim_thompson5910):

(3/4)x or 3x/4

jimthompson5910 (jim_thompson5910):

I'd go with 3x/4

jimthompson5910 (jim_thompson5910):

so the equation is 3x/8 + 240 = 3x/4

OpenStudy (anonymous):

YES I GOT IT! THANK YOU

jimthompson5910 (jim_thompson5910):

you're welcome